OpenStudy (anonymous):

A license plate has to have 3 letters followed by 3 digits. How many license plates can be formed if : 1. the letters and nos. cannot be repeated 2. if the nos. and letters can be repeated. THANKS !!!!!

OpenStudy (anonymous):

2. when nos. and letters are repeated =26*26*26*10*10*10=? 1. when nos. and letters are not repeated =26*25*24*10*9*8=?

OpenStudy (anonymous):

is that right ? yes it seems to be right, I am trying to digest :)